# rQUFguide.bib

### v1

@@ 6 Assumptions and (Counter-) Evidence:
We check the dependence of the LOB state variables to the events’ likelihood of occurrence by plotting the ratio of the joint probability of an event happening with the LOB state variable $v$ i.e. $P(e,v)$ with the product of the individual unconditional probabilities i.e. $P(e)P(v)$, against various categories of the width variable. This ratio informs us whether the two random variables are independent or not.
@@
Queue reactiveness in the LOB has been relatively well established in point process simulations literature ([pakannen2022], [wu2019queue], [huang2015]). It is evident from these papers that the intensities of LOB events depend on the current queue sizes of the LOB. However, it is unclear whether the number of past events alone (i.e., without considering individual quantities) is a reasonable proxy for the cause of this dependence. We would like to note that the Hawkes Process methodology implicitly makes the intensity depend on the count of past events. We perform tests on this assumption in Figure 19. We note that the queue reactiveness is very evident for top of the book queue size in large tick stocks however this reactiveness dies to a significant extent in small tick names. Unsurprisingly, large tick stocks dynamics seem to be unreactive to the deeper quoted volumes but there is a definite structure, particularly in MOs, for small tick stocks.
@@
Clearly from Figure 20(a), we can see that events of the type IS, CO and MO have a definitive trend with increasing $m_{T}$. While in Figure 20(b), there is a U-shape observed in IS and MO order types. Finally in Figure 20(c), we can observe a strong dependence of MO’s probability with increasing $\eta_{IS}$. However we must note that the trend becomes significant only in the right tail of the quantities which are low probability states. We do not see any clear trend when we do the same analysis for $\kappa_{LO},\kappa_{MO}$ (not illustrated). We observe similar characteristics in medium tick stocks and all other small tick stocks as well. Therefore we conclude that though for LOs and COs, we can assume the LOB state variables to be independent of their intensity functions, there exists a clear dependence of intensities of in-spread orders on the current $m_{T},m_{D}$.
Assumption 2.
Order sizes are i.i.d.
Assumption 3.
Order arrival distances from top are i.i.d.
Assumption 4.
Unseen Queue sizes on queue depletion are i.i.d.
To test the i.i.d. assumption, we measure the autocorrelations at various lags of the timeseries of the quantity. It is clear from Figure 18 that Assumption 2 is clearly violated by the empirical data. Assumption 3, however, seems to have some support from the data since the autocorrelations are absent. With regards to Assumption 4, we note that this is a modelling choice.
@@ 7 Discussion and Conclusion
This paper focuses on the disparity between the microstructural behaviour of the LOB between various relative tick-sizes. Tick-sizes not only define the granularity of the price formation process but also impact the market agents’ behaviour. One of our key contributions is to highlight several stylized facts, which are then used to differentiate between large, medium and small tick stocks, and define clear metrics to measure them. We also provide a set of cross-asset visualizations of the stylized facts and show how different (or similar) these attributes are with varying relative tick-size. We make use of a stylized fact known as leverage to model the LOB as a Hawkes Process with the top of the book as a separate dimension and the rest of the book (upto the median of the volume shape of the book) as a separate dimension. We then propose a simple model extending the one proposed in [jain2023hawkes] to account for sparsity, multi-tick level price moves and the humped shape of the book for small-tick stocks. We showcase the universality of the model by performing several simulation studies using randomised parameters in the model. Thereafter, we identify the in-spread Hawkes Parameters and the sparsity variables as the critical variables which define if the simulated LOB will be similar to a large-tick or a small-tick stock. We perform several tests to showcase that a number of stylized facts like sparsity, shape, and relative returns’ distribution can be varied smoothly from that of a large-tick to small-tick asset using our model. We also compare our model’s stylized facts to that of a similar Poisson Process model where we find that the Hawkes model is superior to the Poisson one. Finally, we make note of the assumptions in this simple model and showcase a set of attributes which tests these assumptions. We show that almost all of the model’s assumptions are violated in the real data and conclude by proposing questions for further research in this area in this following.
@@ Acknowledgements
(a) SIRI $Q_{T}$
(b) SIRI $Q_{D}$
(c) AMZN $Q_{T}$
(d) AMZN $Q_{D}$
Figure 19: Testing Assumption 1: Queue Reactiveness: This independence test shows that for large tick stocks like SIRI, the intensity of events like COs, LOs and MOs, and the queue sizes at the top of the book are not independent variables. Unsurprisingly, the same is not true for the deeper volumes in a large tick LOB. For a small tick asset like AMZN, the picture becomes less clear for $Q_{T}$ however for MOs, one can observe a clear dependence among the variables. For $Q_{D}$ too we note that the MOs intensity is quite clearly not independent in the case of small tick stocks.
(a) $m_{T}$
(b) $m_{D}$
(c) $\eta_{IS}$
@@ Appendix A Stylized Facts Visualisations (continued):
In Figure 21, we show the empirical distribution of three stocks: INTC, AAPL, and AMZN, resp. a large, medium and small tick stock. We also report the time-of-day dynamics of the spread noting the ‘tick-constrained’ behaviour in INTC (and other large tick stocks). This is a common term used by practitioners to denote the fact that except for the opening hour, the spread of large tick stocks is constrained to its minimum i.e. one tick.
In Figure 22, we showcase the distance between two successive best quotes for the top 10 quotes in SIRI, PM and BKNG which is resp. a large, medium and small tick stock. We conclude that distribution of empty levels is similar across all depths and it becomes flatter as we decrease the rel. tick-size.
@@
(a) SIRI (Large tick)
(b) PM (Medium tick)
(c) BKNG (Small tick)
@@
(a) IBM (Medium tick)
(b) AMZN (Small tick)
Figure 24: Leverage (from top): Darker the hue, more the leverage
@@
Figure 25, we show per stock year wide average shapes (in red) as well as daily average shapes (in light blue). Firstly the shape is stable across the whole dataset i.e. for the whole year. Secondly, we again stress the point made in Section 3.3 that the shape transforms from a decaying shape for large and medium tick stocks to that where the maxima is actually deeper in the book for small tick stocks. The distribution of this maxima’s instantaneous value is also shown in Figure 26.
Appendix B Simulation Study Results (continued):

### v2

@@ 5.2 Assumptions and Their Empirical Evidence or Counter-Evidence
(a) SIRI $Q_{T}$
(b) AMZN $Q_{T}$
Figure 18: Testing Assumption 1: Queue Reactiveness: This independence test shows that for large tick stocks like SIRI, the intensity of events like COs, LOs and MOs, and the queue sizes at the top of the book are not independent variables. Unsurprisingly, the same is not true for the deeper volumes in a large tick LOB. For a small tick asset like AMZN, the picture becomes less clear for $Q_{T}$ however for MOs, one can observe a clear dependence among the variables. For $Q_{D}$ too we note that the MOs intensity is quite clearly not independent in the case of small tick stocks (check Appendix Figure 28).
@@
We check the dependence of the LOB state variables to the events’ likelihood of occurrence by plotting the ratio of the joint probability of an event happening with the LOB state variable $v$ i.e. $P(e,v)$ with the product of the individual unconditional probabilities i.e. $P(e)P(v)$, against various categories of the width variable. This ratio informs us whether the two random variables are independent or not. While this test is not exhaustive or strict, it does inform us of the interdependence of two random variables in a simplified manner. If the ratio is near one, we have some albeit insufficient evidence to consider the two variables independent. However if the ratio is significantly different from one, we can conclude that the two variables are not independent.
@@
Queue reactiveness in the LOB, i.e. the notion that event intensities depend directly on the current sizes of the queues in the LOB ($Q^{(.)}_{(.)}(t)$), has been relatively well established in point process simulations literature ([pakannen2022], [wu2019queue], [huang2015]). It is evident from these papers that the intensities of LOB events depend on the current queue sizes of the LOB. However, it is unclear whether the number of past events alone (i.e., without considering individual quantities) is a reasonable proxy for the cause of this dependence. We would like to note that the Hawkes Process methodology implicitly makes the intensity depend on the count of past events. We perform tests on this assumption in Figure 18. We note that the queue reactiveness is very evident for top of the book queue size in large tick stocks however this reactiveness dies to a significant extent in small tick names.
@@
Further we show the same analysis for $Q_{D}$ instead of $Q_{T}$ in Figure 28. Unsurprisingly, large tick stocks dynamics seem to be unreactive to the deeper quoted volumes but there is a definite structure, particularly in MOs, for small tick stocks. Additionally, we see that events of the type IS, CO and MO have a definitive trend with increasing $m_{T}$ (Figure 29(a) in the Appendix). While there is a U-shape observed in IS and MO order types with increasing $m_{D}$ (see Appendix Figure 29(b)). Finally we observe a strong dependence of MO’s probability with increasing $\eta_{IS}$ (in the Appendix Figure 29(c)). However we must note that the trend becomes significant only in the right tail of the quantities which are low probability states. We do not see any clear trend when we do the same analysis for $\kappa_{LO},\kappa_{MO}$ (not illustrated). We observe similar characteristics in medium tick stocks and all other small tick stocks as well. Therefore we conclude that though for LOs and COs, we can assume the LOB state variables to be independent of their intensity functions, there exists a clear dependence of intensities of in-spread orders on the current $m_{T},m_{D}$.
@@
Assumption 2.
(a) Order sizes, (b) Order arrival distances and (c) Unseen Queue sizes on queue depletion, are i.i.d. random variables.
To test the i.i.d. assumption, we measure the autocorrelations at various lags of the timeseries of the quantity. It is clear from Figure 19 that Assumption 2a is clearly violated by the empirical data. Assumption 2b, however, seems to have some support from the data since the autocorrelations are absent. With regards to Assumption 2c, we note that this is a modelling choice.
@@ 6 Discussion and Conclusion
In this work we outline how, statistically, does the LOB change when the granularity of price levels is kept constant with a constant tick size while the price of the security increases from a few dollars to a few thousands of dollars. We quantify this change by the change in relative tick size which is defined as the ratio between the tick size and the price of the security. One of our key contributions is to highlight several stylized facts, which are then used to differentiate between large, medium and small tick stocks, and define clear metrics to measure them. We also provide a set of cross-asset visualizations of the stylized facts and show how different (or similar) these attributes are with varying relative tick-size. We found the following power law relations in these stylized facts. Firstly, the average spread of a security and the average absolute returns of the security are both approximately proportional to the inverse of the relative tick-size. Secondly, the depth at which we find the maximum liquidity in the LOB follows a power law over the relative tick-size with the exponent roughly equal to $\frac{3}{2}$. In addition to that we see definitive trends in the shape of the spreads distribution as well as the sparsity in the LOB as well decrease the relative tick-size. We make use of a stylized fact known as leverage, which seems to be universally similar in structure across all stocks we studied, to model the LOB as a Hawkes Process with the top of the book as a separate dimension and the rest of the book (upto the median of the volume shape of the book) as a separate dimension.
We propose a simple model extending the one proposed in [jain2023hawkes] to account for sparsity, multi-tick level price moves and the humped shape of the book for small-tick stocks. We showcase the universality of the model by performing several simulation studies using randomised parameters in the model. Thereafter, we identify the in-spread Hawkes Parameters and the sparsity variables as the critical variables which define if the simulated LOB will be similar to a large-tick or a small-tick stock. We perform several tests to showcase that a number of stylized facts like sparsity, shape, and relative returns’ distribution can be varied smoothly from that of a large-tick to small-tick asset using our model. We also compare our model’s stylized facts to that of a similar Poisson Process model where we find that the Hawkes model is superior to the Poisson one.
Finally, we make note of the assumptions in this simple model and showcase a set of attributes which tests these assumptions. We show that almost all of the model’s assumptions are violated in the real data and conclude by proposing questions for further research in this area in this following.
@@ Appendix A Stylized Facts Visualisations (continued):
In Figure 20, we show the empirical distribution of three stocks: INTC, AAPL, and AMZN, resp. a large, medium and small tick stock. We also report the time-of-day dynamics of the spread noting the ‘tick-constrained’ behaviour in INTC (and other large tick stocks). This is a common term used by practitioners to denote the fact that except for the opening hour, the spread of large tick stocks is constrained to its minimum i.e. one tick. In the time-of-day analysis, we observe tick-constrained dynamics in large-tick stocks like INTC, where the spread remains floored at one tick throughout the trading day, despite price discovery causing fluctuations in the spread. This behavior is markedly different in medium and small-tick stocks, where the spread exhibits more variation throughout the day.
@@
The shape of the book decays from its maximum at half a tick from the mid-price for large and medium-tick stocks, as illustrated in Figures 22(a) and 22(b). Notably, the shape profile decays more rapidly for large-tick stocks compared to medium-tick ones. For stocks like IBM, which are close to the boundary between medium and small-tick stocks, the maximum liquidity shifts deeper into the book. On the other hand, for small-tick stocks such as the one shown in Figure 22(c), the liquidity distribution is wider, with the maximum liquidity lying several ticks below the mid-price.
The empirical density of the maxima of the liquidity as a function of distance from the mid-price, shown in Figure 23, indicates that the variance in the location of maximum liquidity increases as the relative tick size decreases. The variability in the location of this maximum provides insights into how deep the liquidity is distributed in the book. Even in small-tick stocks, there are significant instances where the maximum liquidity lies just half a tick from the mid-price. However, as depicted in Figure 20, the time-weighted likelihood of a one-tick wide spread is quite low.
@@
(a) SIRI (Large tick)
(b) BKNG (Small tick)
In Figure 24, we showcase the distance between two successive best quotes for the top 10 quotes in SIRI, PM and BKNG which is resp. a large, medium and small tick stock. We conclude that distribution of empty levels is similar across all depths and it becomes flatter as we decrease the rel. tick-size. Next, we calculate the percentiles of the shape of the book at the 10th, 25th, 75th, and 90th percentiles. The conjecture is that these percentiles will remain relatively invariant if the book is dense, meaning there are close to zero empty levels in the LOB. Thus, the variance of these percentile levels serves as a proxy measurement of the LOB’s sparsity. As shown in Figure 25, there is a steady relationship between the variance of the levels where certain percentages of total liquidity can be found and the relative tick size. This variance increases with decreasing tick size, indicating greater sparsity.
Figure 25: Variance of Volume Percentiles (log)
In Figure 25, we show the variability in the instantaneous shape of the LOB by calculating the variance of the percentiles of the shape of the LOB. This measure informs us about the variation of liquidity in the book and therefore is a proxy measure for sparsity. As we can see in this figure, there is continuous trend of increasing variance as we decrease the rel. tick size for all percentiles of the shape of the LOB.
(a) IBM (Medium tick)
(b) AMZN (Small tick)
Figure 26: Leverage (from top): Darker the hue, more the leverage
@@
Appendix B Simulation Study Parameters (continued):
@@
Appendix C Assumptions and Counter-Evidence (continued):
(a) SIRI $Q_{D}$
(b) AMZN $Q_{D}$
Figure 28: Testing Assumption 1: Queue Reactiveness: For $Q_{D}$ too we note that the MOs intensity is quite clearly not independent in the case of small tick stocks.
(a) $m_{T}$
(b) $m_{D}$
(c) $\eta_{IS}$